.Gallatin Carpet Cleaning is a small, family-owned business operating out of Bozeman, Montana. For its services, the company hResource consumption is distributed across the activities as follows: Job Support Other 0% Distribution of Resource ConsumptiRequired 1 Required 2 Required 3 Required 4 Prepare the first stage allocation of costs to the activity cost pools. Cleaning4. The revenue from the Flying N Ranch was $138.00 (600 square feet @ $23.00 per hundred square feet). Calculate the customerComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 The revenue frComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 The company re
Answer #1
1
Cleaning Carpets Travel to Jobs Job support Other Total
Wages 111390 18330 0 11280 141000
Cleaning supplies 33000 0 0 0 33000
Cleaning equipment depreciation 7260 0 0 3740 11000
Vehicle expenses 0 24650 0 4350 29000
Office expense 0 0 34220 24780 59000
President’s compensation 0 0 24140 46860 71000
Total cost 151650 42980 58360 91010 344000
2
Activity Cost Pool Activity Rate   
Cleaning carpets 17.84 per hundred square feet =151650/8500
Travel to jobs 0.62 per mile =42980/69000
Job support 29.18 per job =58360/2000
3
Activity Cost Pool Activity Rate    Activity Total
Cleaning carpets 17.84 6 107.04
Travel to jobs 0.62 60 37.20
Job support 29.18 1 29.18
173.42
Cost of the job 173.42
4
Customer Margin—Activity-Based Costing
Sales 138.00
Costs:
Cleaning carpets 107.04
Travel to jobs 37.20
Job support 29.18 173.42
Customer margin -35.42